# Design Stability in Adaptive Experiments: Implications for Treatment Effect Estimation

### v1

@@ Abstract
We study the problem of estimating the average treatment effect (ATE) under sequentially adaptive treatment assignment mechanisms. In contrast to classical completely randomized designs, we consider a setting in which the probability of assigning treatment to each experimental unit may depend on prior assignments and observed outcomes. Within the potential outcomes framework [1], we propose and analyze two natural estimators for the ATE: the inverse propensity weighted (IPW) estimator and an augmented IPW (AIPW) estimator. The cornerstone of our analysis is the concept of design stability, which requires that as the number of units grows, either the assignment probabilities converge, or sample averages of the inverse propensity scores and of the inverse complement propensity scores converge in probability to fixed, non-random limits. Our main results establish central limit theorems for both the IPW and AIPW estimators under design stability and provide explicit expressions for their asymptotic variances. We further propose estimators for these variances, enabling the construction of asymptotically valid confidence intervals. Finally, we illustrate our theoretical results in the context of Wei’s adaptive coin design [2] and Efron’s biased coin design [3], highlighting the applicability of the proposed methods to sequential experimentation with adaptive randomization.
@@ 1 Introduction
Estimating the average treatment effect is a foundational problem in causal inference, especially when evaluating interventions in fields such as healthcare [4], education [5], public policy [6], development economics [7], and digital experimentation [8]. Traditional methods often assume simple randomized designs with independent and identically distributed (i.i.d.) units and fixed treatment assignment probabilities. However, many real-world experiments depart from this idealized setting: units often arrive sequentially, treatment assignments may adapt over time based on previous allocations, observed outcomes, or covariate information, and the study population is finite. Such sequential finite-population setups are common in adaptive clinical trials, online A/B testing, and policy evaluations. In these scenarios, adaptively assigning treatments can lead to complex dependencies among units, causing traditional ATE estimators to become biased or inefficient and undermining the applicability of conventional asymptotic results. Despite its practical importance, this setting remains methodologically less explored.
@@
This paper develops a general framework for ATE estimation and inference under sequential designs in finite populations. We study the asymptotic behavior of the average treatment effect estimators under a class of sequential Bernoulli assignment mechanisms, in which the probability of assigning treatment to unit $i$ may depend on the observed history up to that point. Formally, $\mathbb{P}(K_{i}=1\mid\mathcal{F}_{i-1})=p_{i},$ $p_{i}\in\mathcal{F}_{i-1},$ where $K_{i}=1$ indicates that the $i$th unit is assigned to treatment, and $\mathcal{F}_{i-1}$ denotes the sigma-field generated by the treatment assignments and outcomes of the first (i1)(i-1) units. Two classic examples of such assignments that will be discussed in this paper are Wei’s adaptive coin design [2], and Efron’s biased coin design [3]. In the former, the assignment probabilities are expressed as a non-increasing function of the relative imbalance until the previous step, whereas in the latter design these probabilities take constant values that depend on the treatment-control imbalance up to the previous step (for example, fixed values like $\eta$, (1-$\eta$), $\tfrac{1}{2}\bigr)$.
@@
Our contributions proceed in four parts. First, we begin by analyzing a standard estimator of the average treatment effect: the inverse propensity weighting (IPW) estimator, and propose an improvement by introducing a finite-population version of its augmented version (AIPW) that is commonly defined and used in model-based frameworks. Second, for both estimators, we establish central limit theorems under general sequential designs that satisfy a newly defined property called design stability. Third, under two different forms of design stability - strong and weak - we derive estimators of the asymptotic variances of the treatment effect estimators. These variance estimators, and the corresponding confidence intervals for the ATE, are conservative in that they are asymptotically positively biased, leading to asymptotic overcoverage of the confidence intervals. However, the biases vanish under certain forms of treatment effect homogeneity, yielding correct asymptotic coverage of the confidence intervals. Finally, we specialize these results to the two concrete experimental designs mentioned above, arguing that one of them (Wei’s adaptive design [2]) satisfies the strong design stability condition, whereas the other (Efron’s design [3]) satisfies the weak design stability condition.
@@
The remainder of the paper is organized as follows. Section 2 reviews relevant prior work. Section 3 formally defines the problem, introduces the potential outcomes framework, describes the sequential assignment structure, and presents the estimators of interest. Section 4 presents the main theoretical results, including central limit theorems for adaptive designs and conservative asymptotic variance estimators for confidence interval construction. These results are then specialized in Section 5 to two widely used adaptive treatment assignment mechanisms. Sections 5.1 and 5.2 examine the stability of these designs and present simulation studies illustrating the finite-sample performance of the proposed estimators and supporting the theoretical findings. Section 6 concludes with a discussion and directions for future research. Proofs of all main and auxiliary results are provided in the supplementary material.
@@ 2 Related work
Causal inference in experimental settings is typically framed through two paradigms. The infinite-superpopulation perspective views study units as random draws from an underlying population, with randomness arising from the data-generating process. In contrast, the finite-population or design-based perspective treats the set of units as fixed, with uncertainty introduced solely through the experimental design. This latter view, combined with the potential outcomes formulation, traces back to [1], who conceptualized each unit’s treatment and control responses as fixed quantities and attributed randomness entirely to randomization. While classical asymptotic theory [[, e.g.,]]lehmann1999, vandervaart2000 is often aligned with the superpopulation framework, many applications, particularly randomized trials and survey sampling, are more naturally analyzed from the finite-population perspective [[, e.g.,]]fisher1935, cochran1977, rosenbaum2002, imbens2015.
@@
The study of asymptotic normality in causal inference can be traced back to results on simple random sampling. Classical central limit theorems were established by [15][16][17], with convenient formulations presented in [18, 9]. These sampling-based central limit theorems can also be viewed as special cases of the more general results for rank statistics [19, 20, 21, 22]. Further foundational work includes the theory of U-statistics developed by [23] and the weak convergence results of [24], which laid the groundwork for modern asymptotic theory in survey sampling and experimental design. Because treatment and control groups in randomized experiments correspond to simple random samples from the finite set of experimental units, these sampling-based central limit theorems are directly applicable to the difference-in-means estimator of the average treatment effect. This connection underlies much of the early asymptotic justification in randomization-based causal inference [[, e.g.,]]Liu2014-kk, Ding2018-lr, Fogarty2018, li2017general.
@@
In modern applications such as adaptive clinical trials [29], online A/B tests [30], and adaptive policy experiments [31], treatment assignments may evolve in response to interim data, violating the independence assumptions of static designs. Such sequential mechanisms introduce dependence across units, requiring martingale‐based CLTs [32, Chapter 3] in place of classical i.i.d. arguments.
In spite of the recent explosion of research on design-based finite population inference, to the best of our knowledge, rigorous theory for finite-population central limit theorems under general sequential general sequential Bernoulli assignments remains scarce. Recent explorations on inference of ATE from adaptive designs have been done in a setting where the potential outcomes for each experimental unit are assumed to follow an unknown probability distribution $\mathcal{P}$ and the ATE is defined as the difference of expectations of the potential outcomes with respect to $\mathcal{P}$. In this setting, [33] established asymptotic normality of the difference-in-means estimator under an adaptive Bernoulli allocation rule, and [34] extended these results to the augmented inverse probability weighting (AIPW) estimator [35]. However, this setting is different from design or randomization-based inference, where the uncertainty in the data (and consequently in the estimator) is induced solely by the act of randomization.
@@
The present work addresses this gap, establishing central limit theorems for IPW and AIPW-inspired estimators in finite populations under broad sequential designs in a purely design-based inferential framework, where the potential outcomes are assumed fixed.
@@ 3 Problem Description
Consider a study with $N$ experimental units indexed by $i=1,\ldots,N$. We adopt the potential outcomes framework, introduced by [1] and later formalized by [36]. For each unit $i$, the outcome of interest $Y_{i}$ is characterized by two potential outcomes: $Y_{i}(0)$ under control and $Y_{i}(1)$ under treatment. The individual treatment effect is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and our target parameter is the average treatment effect (ATE), defined as
@@
for some constant $\tau\in\mathbb{R}$ is called additivity of potential outcomes and is standard in literature [[, e.g.,]Chapter 6]imbens2015. Here we introduce the following definition that generalizes the concept of treatment effect homogeneity:
@@
$Y_{i}(1)-\overline{Y}_{N}(1)\propto Y_{i}(0)-\overline{Y}_{N}(0)\quad\text{for all }i,$ (3)
where $\overline{Y}_{N}(\ell)=\tfrac{1}{N}\sum_{i=1}^{N}Y_{i}(\ell)$ for $\ell\in\{0,1\}$.
@@
for some constant $c\in\mathbb{R}$. We will see that conditions (2)-(4) play important roles in the inference problem to be discussed.
In the classical randomized treatment allocation design, a pre-defined constant number $N_{1}$ of the $N$ units are assigned to treatment, with the subset selected uniformly at random [36]. Formally, let $\mathbf{K}=(K_{1},K_{2},\ldots,K_{N})^{\text{T}}\in\{0,1\}^{N}$ denote the random assignment vector, where $K_{i}=1$ if $i$th unit is assigned to the treatment group and $K_{i}=0$ otherwise. A simple random sample of size $N_{1}$ is chosen from the finite population using the assignment vector $\mathbf{K}$, where $\mathbb{P}(\mathbf{K}=\mathbf{k})=\binom{N}{N_{1}}^{-1}$ for all $\mathbf{k}\in\{0,1\}^{N}$ satisfying $\mathbf{1}_{N}^{\text{T}}\mathbf{k}=N_{1}$. Given a treatment assignment vector $\mathbf{k}\in\{0,1\}^{N}$, the observed data $\{Y_{i}\}_{i=1}^{N}$ are the realized potential outcomes, where each unit’s outcome corresponds to its assigned treatment or control, defined by
@@
where $N_{1}=\sum_{i=1}^{N}K_{i}\text{ and }N_{0}=N-N_{1}.$ This estimator is unbiased and satisfies a central limit theorem as $N$ grows to infinity [28]. We note that the difference-in-means estimator is a special case of the Horvitz–Thompson type estimator [37], also known as the inverse propensity weighted (IPW) estimator, defined as
$\widehat{\tau}_{\textsc{IPW}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left\{\frac{K_{i}Y_{i}}{p_{i}}-\frac{(1-K_{i})Y_{i}}{1-p_{i}}\right\},$ (6)
where the weights $p_{i}=N_{1}/N$ for $i=1,\ldots,N$.
In contrast to the classical randomized treatment assignment, we consider in this paper a sequential treatment assignment mechanism, in which the probability of assigning the $i$th unit to treatment is adaptive. In other words, for each unit $i\in{1,2,\ldots,N}$, the assignment indicator $K_{i}$, conditional on the past history, follows a Bernoulli distribution with success probability $p_{i}$, which we refer to as the inclusion probability. The inclusion probability $p_{i}$ is not fixed; rather, it is a measurable function of the prior assignment history and outcomes. Formally, we define a sequence of increasing sigma-fields $\{\mathcal{F}_{i-1}\}_{i\geq 1}$, where $\mathcal{F}_{i-1}=\sigma(K_{1},Y_{1},\ldots,K_{i-1},Y_{i-1})$ represents the cumulative information available after assigning treatment or control and observing the outcome of the (i1)(i-1)th unit. The assignment mechanism is such that
@@
In words, the probability of assigning treatment to unit $i$ may depend on all previous assignments and observations up to stage (i1)(i-1) in an arbitrary way; we call this assignment a sequential treatment assignment. As an example, one may choose $p_{i}$ to promote relative balance between the numbers of treatment and control assignments. In this case, $p_{i}$ can be defined as the complement of the moving average of past assignments:
@@
In this assignment, if the previous units have mostly been assigned treatment, the chance of assigning treatment to the next unit will be lowered and vice versa [2].
@@
It is well known that IPW estimators suffer from inflated variance when the probabilities approach extremes [38]. In a model-based setting, this limitation of IPW estimator is mitigated by the AIPW estimator [35] via model augmentation, offering double robustness. In our setting, where no probability model for the potential outcomes is assumed, we propose the following finite-population model-free version of the AIPW estimator:
@@
Note that for large NN, the weighted average $\widehat{Y}_{N}(\ell)$ serves as an intuitive estimator of $\overline{Y}_{N}(\ell)$ for $\ell\in\{0,1\}$. In this sense, $\widehat{\tau}_{\textsc{AIPW}}$ is directly motivated from the classical AIPW estimator [35]. We formalize this intuition in a later theorem on the behavior of the AIPW estimator $\widehat{\tau}_{\textsc{AIPW}}$ (see Theorem 4).

### v2

@@ Abstract
We study the problem of estimating the average treatment effect (ATE) under sequentially adaptive treatment assignment mechanisms. In contrast to classical completely randomized designs, we consider a setting in which the probability of assigning treatment to each experimental unit may depend on prior assignments and observed outcomes. Within the potential outcomes framework [Neyman1923], we propose and analyze two natural estimators for the ATE: the inverse propensity weighted (IPW) estimator and an augmented IPW (AIPW) estimator. The cornerstone of our analysis is the concept of design stability, which requires that as the number of units grows, either the assignment probabilities converge, or sample averages of the inverse propensity scores and of the inverse complement propensity scores converge in probability to fixed, non-random limits. Our main results establish central limit theorems for both the IPW and AIPW estimators under design stability and provide explicit expressions for their asymptotic variances. We further propose estimators for these variances, enabling the construction of asymptotically valid confidence intervals. Finally, we illustrate our theoretical results in the context of Wei’s adaptive coin design [Wei1978-gd] and Efron’s biased coin design [Efron1971-yh], highlighting the applicability of our results to sequential experimental designs with adaptive randomization.
@@ 1 Introduction
Estimating the average treatment effect is a foundational problem in causal inference, especially when evaluating interventions in fields such as healthcare [Shen2020-zr], education [Kaplan2016-ok], public policy [Finkelstein2020-jw], development economics [Imbens2004-nc], and digital experimentation [Shi2023-at]. Traditional methods often assume simple randomized designs with independent and identically distributed (i.i.d.) units and fixed treatment assignment probabilities. However, many real-world experiments depart from this idealized setting: units often arrive sequentially, treatment assignments may adapt over time based on previous allocations, observed outcomes, or covariate information, and the study population is finite. Such sequential finite-population setups are common in adaptive clinical trials, online A/B testing, and policy evaluations. In these scenarios, adaptively assigning treatments can lead to complex dependencies among units, causing traditional ATE estimators to become biased or inefficient and undermining the applicability of conventional asymptotic results. Despite its practical importance, this setting remains methodologically less explored.
@@
This paper develops a general framework for ATE estimation and inference under sequential designs in finite populations. We study the asymptotic behavior of the average treatment effect estimators under a class of sequential Bernoulli assignment mechanisms, in which the probability of assigning treatment to unit $i$ may depend on the observed history up to that point. Two classic examples of such assignments are Wei’s adaptive coin design [Wei1978-gd], and Efron’s biased coin design [Efron1971-yh]. More general classes of adaptive assignments in clinical trials can be found in [bb691c50-25c9-325f-b1ee-0cc92167cbd1].
@@
Our contributions proceed in four parts. First, we analyze the inverse propensity weighted (IPW) estimator for the average treatment effect and propose an improvement via a finite-population analogue of its augmented version (AIPW), commonly used in model-based frameworks. Second, for both estimators, we establish central limit theorems under general sequential designs that satisfy a newly defined property called design stability. Third, under two different forms of design stability - strong and weak - we derive estimators of the asymptotic variances of the average treatment effect estimators. These variance estimators, and the corresponding confidence intervals for the ATE, are conservative in that they are asymptotically positively biased, leading to asymptotic overcoverage of the confidence intervals. However, the biases vanish under certain forms of treatment effect homogeneity, yielding correct asymptotic coverage of the confidence intervals. Finally, we specialize these results to the two concrete experimental designs mentioned above, arguing that one of them (Wei’s adaptive coin design [Wei1978-gd]) satisfies the strong design stability condition, whereas the other (Efron’s biased coin design [Efron1971-yh]) satisfies the weak design stability condition.
@@
The remainder of the paper is organized as follows. Section 2 reviews relevant prior work. Section 3 formally defines the problem, introduces the potential outcomes framework, describes the sequential assignment structure, and presents the estimators of interest. Section 4 presents the main theoretical results, including central limit theorems for adaptive designs and conservative asymptotic variance estimators for confidence interval construction. These results are then specialized in Section 5 to two widely used adaptive treatment assignment mechanisms. Sections 5.1 and 5.2 examine the stability of these designs and present simulation studies illustrating the finite-sample performance of the proposed estimators and supporting the theoretical findings. Section 5.3 briefly discusses extensions to more general adaptive clinical trial designs, while Section 6 concludes with a discussion and directions for future research. Proofs of all main and auxiliary results are provided in the Supplementary Material.
@@ 2 Related work
Causal inference in experimental settings is typically framed within two paradigms. The infinite superpopulation perspective views study units as random draws from an underlying population, with randomness arising from the data-generating process. In contrast, the finite-population or design-based perspective treats the set of units as fixed, with uncertainty introduced solely through the experimental design. This latter view, combined with the potential outcomes formulation, traces back to [Neyman1923], who conceptualized each unit’s treatment and control responses as fixed quantities and attributed randomness entirely to randomization. While classical asymptotic theory [[, e.g.,]]lehmann1999, vandervaart2000 is often aligned with the superpopulation framework, many applications, particularly randomized trials and survey sampling, are more naturally analyzed from the finite-population perspective [[, e.g.,]]fisher1935, cochran1977, rosenbaum2002, imbens2015.
@@
The study of asymptotic normality in causal inference can be traced back to results on simple random sampling. Classical central limit theorems were established by [madow1948][erdos1959][hajek1960], with convenient formulations presented in [lehmann1975, lehmann1999]. These sampling-based central limit theorems can also be viewed as special cases of the more general results for rank statistics [wald1944, Noether1949-br, Fraser1956-qf, hajek1961]. Further foundational work includes the theory of U-statistics developed by [Hoeffding1963-tc] and the weak convergence results of [1177692698], which laid the groundwork for modern asymptotic theory in survey sampling and experimental design. Because treatment and control groups in randomized experiments correspond to simple random samples from the finite set of experimental units, these sampling-based central limit theorems are directly applicable to the difference-in-means estimator of the average treatment effect. This connection underlies much of the early asymptotic justification in randomization-based causal inference [[, e.g.,]]Liu2014-kk, Ding2018-lr, Fogarty2018, li2017general.
@@
In modern applications such as adaptive clinical trials [rosenberger2016, Hu2006-vc, bb691c50-25c9-325f-b1ee-0cc92167cbd1], online A/B tests [article], and adaptive policy experiments [ATHEY201773], treatment assignments may evolve in response to interim data, violating the independence assumptions of static designs. Such sequential mechanisms introduce dependence across units, requiring martingale‐based techniques [Hall1980-ny, Chapter 3] in place of classical i.i.d. arguments.
There is a growing literature on inference under adaptive treatment assignment and bandit-style designs adopts stability-type assumptions and develops martingale-based central limit theorems, leading to methodological parallels with our work; examples include [Kato2020-qo], [pmlr-v236-cook24a], [NEURIPS2021_eff30581], [Hadad2021-jg], and [NEURIPS2020_6fd86e0a, Zhang2021-zy]. There are, however, several fundamental differences. First, all of these works operate under a superpopulation framework, where randomness in observed outcomes arises from assumed probabilistic outcome models, whereas we adopt a finite-population perspective in which the sole source of randomness is the known probabilistic treatment assignment mechanism. This distinction leads to substantively different assumptions and technical arguments. Second, while some of the above papers ([Hadad2021-jg]; [Zhang2021-zy]) explicitly use a potential outcomes framework, others do not. Third, despite superficial similarities in the notion of “stability,” the bandit algorithms studied in [NEURIPS2020_6fd86e0a, Zhang2021-zy] are designed to eventually commit to a single arm, in sharp contrast to our setting, where long-run balance across treatment arms, which is central to randomized experimentation, is the primary objective.
@@
The present work addresses this gap, establishing central limit theorems for IPW and AIPW-inspired estimators in finite-population under broad sequential designs in a purely design-based inferential framework, where the potential outcomes are assumed fixed.
@@ 3 Problem Description
Consider a study with $N$ experimental units, indexed by $i=1,\ldots,N$. We adopt the potential outcomes framework, introduced by [Neyman1923] and later formalized by [Rubin1974-gr]. For the $i$th unit, the outcome of interest $Y_{i}$ is characterized by two potential outcomes: $Y_{i}(0)$ under control and $Y_{i}(1)$ under treatment. The individual treatment effect is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and our target parameter is the average treatment effect (ATE), defined as
@@
for some constant $\tau\in\mathbb{R}$, is referred to as additivity of potential outcomes and is standard in literature [[, e.g.,]Chapter 6]imbens2015. Here we introduce the following definition that generalizes the concept of treatment effect homogeneity.
@@
$Y_{i}(1)-\overline{Y}_{N}(1)=\lambda(Y_{i}(0)-\overline{Y}_{N}(0))\quad\text{for all }i,$ (3)
for some constant $\lambda\in\mathbb{R}$, where $\overline{Y}_{N}(\ell)=\tfrac{1}{N}\sum_{i=1}^{N}Y_{i}(\ell)$ for $\ell\in\{0,1\}$.
@@
for some constant $c\in\mathbb{R}$. We will see that the conditions (2)-(4) play important roles in the inference problem to be discussed.
In the classical randomized treatment allocation design, a pre-defined constant number $N_{1}$ of the $N$ units are assigned to treatment, with the subset selected uniformly at random [Rubin1974-gr]. Formally, let $\mathbf{K}=(K_{1},K_{2},\ldots,K_{N})^{\text{T}}\in\{0,1\}^{N}$ denote the random assignment vector, where $K_{i}=1$ if $i$th unit is assigned to the treatment group and $K_{i}=0$ otherwise. A simple random sample of size $N_{1}$ is chosen from the finite population using the assignment vector $\mathbf{K}$, where $\mathbb{P}(\mathbf{K}=\mathbf{k})=\binom{N}{N_{1}}^{-1}$ for all $\mathbf{k}\in\{0,1\}^{N}$ satisfying $\mathbf{1}_{N}^{\text{T}}\mathbf{k}=N_{1}$. Given a treatment assignment vector $\mathbf{k}\in\{0,1\}^{N}$, the observed data $\{Y_{i}\}_{i=1}^{N}$ are the realized potential outcomes, where each unit’s outcome corresponds to its assigned treatment or control, defined by
@@
where $N_{1}=\sum_{i=1}^{N}K_{i}\text{ and }N_{0}=N-N_{1}.$ This estimator is unbiased and satisfies a central limit theorem as $N$ grows to infinity [li2017general]. We note that the difference-in-means estimator is a special case of the Horvitz–Thompson type estimator [horvitz1952], also known as the inverse propensity weighted (IPW) estimator, defined as
$\widehat{\tau}_{\textsc{IPW}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left\{\frac{K_{i}Y_{i}}{p_{i}}-\frac{(1-K_{i})Y_{i}}{1-p_{i}}\right\}.$ (6)
In particular, when $p_{i}=N_{1}/N$ for all $i$, this estimator reduces to the usual difference-in-means estimator.
In contrast to the classical randomized treatment assignment, we consider in this paper a sequential treatment assignment mechanism, in which the probability of assigning the $i$th unit to treatment is adaptive. In other words, for each unit $i\in\{1,2,\ldots,N\}$, the assignment indicator $K_{i}$, conditional on the past history, follows a Bernoulli distribution with success probability $p_{i}$, which we refer to as the inclusion probability. The inclusion probability $p_{i}$ is not fixed; rather, it is a measurable function of the prior assignment history and outcomes. Formally, we define a sequence of increasing sigma-fields $\{\mathcal{F}_{i-1}\}_{i\geq 1}$, where $\mathcal{F}_{i-1}=\sigma(K_{1},Y_{1},\ldots,K_{i-1},Y_{i-1})$ represents the cumulative information available after assigning treatment or control and observing the outcome of the (i1)(i-1)th unit. The assignment mechanism is such that
@@
In words, the probability of assigning treatment to the $i$th unit may depend on all previous assignments and observations up to stage (i1)(i-1) in an arbitrary way; we call this assignment a sequential treatment assignment. As an example, one may choose $p_{i}$ to promote relative balance between the numbers of treatment and control assignments. In this case, $p_{i}$ can be defined as the complement of the moving average of past assignments:
@@
In this assignment, if the previous units have mostly been assigned treatment, the chance of assigning treatment to the next unit will be lowered and vice versa [Wei1978-gd].
@@
It is well known that IPW estimators suffer from inflated variance when the probabilities approach extremes [tsiatis2006]. In a model-based setting, this limitation of IPW estimator is mitigated by the AIPW estimator [robins1994] via model augmentation, offering double robustness. In our setting, where no probability model for the potential outcomes is assumed, we propose the following finite-population model-free version of the AIPW estimator:
@@
Note that for large $N$, the weighted average $\widehat{Y}_{N}(\ell)$ serves as an intuitive estimator of $\overline{Y}_{N}(\ell)$ for $\ell\in\{0,1\}$. In this sense, $\widehat{\tau}_{\textsc{AIPW}}$ is directly motivated from the classical AIPW estimator [robins1994]. We formalize this intuition in a later theorem on the behavior of the AIPW estimator $\widehat{\tau}_{\textsc{AIPW}}$ (see Theorem 4).